OpenStudy (omgthatgirl):

Solve by factoring or finding square roots X^2-3x-4=0 A. 2, -2 B. -4. -1 C. 2, -1 D. 4, -1

OpenStudy (anonymous):

(x + 4 ) (x - 1) check by foiling: First: x * x = x^2 Outter: x * -1= -x Inner: 4 * x = 4x Last: 4 * -1= -4 add like terms: 4x -x = 3x so it works now set binomials to equal 0: x + 4 = 0 subtract 4 x= -4 x -1 =0 add 1 x = 1 x= -4, 1